OpenStudy (rissaunnie):

MEDAL AND FAN IMPORTANT graph the linear equation by plotting three points -3y=-2x-6

OpenStudy (phi):

to make the equation a bit simpler I would divide both sides by -3 \[ \frac{-3y}{-3} = \frac{-2x-6}{-3} \] that becomes \[ y = \frac{-2x}{-3} + \frac{-6}{-3} \\ y= \frac{2}{3}x + 2 \]

OpenStudy (phi):

next, pick easy numbers for x I would use x=-3 x=0 and x=3 do you know how to find y ?

OpenStudy (sushi121212):

 you have to divide the sides by 3 -3y=-2x-6 __ ____ -3 -3 That would get you: y=2/3x+2